For the current month, the owner of the workshop received an income of UAH 200 thousand. From this income, he must pay 5 thousand UAH. to his assistant and UAH 2.5 thousand. cleaning lady. The cost of raw materials and materials spent on production is 170 thousand UAH.

Determine the accounting and net economic profit if the owner of the workshop could receive a salary of 7.5 thousand UAH. and a return on investment of 15% per annum. (When organizing the enterprise, the owner invested 100 thousand UAH in it.)

 



1
Expert's answer
2021-02-25T18:17:18-0500
"TR=200000"

"TC=5000+2500+170000=177500"

accounting profit


"\pi=TR-TC=22500"

economic profit


"22500-7500-\frac {0.15\times100000}{12}=13750"